OpenStudy (anonymous):

Will give medal. Use cross products to demonstrate whether or not the two ratios are equivalent: 4/6 and 14/21

OpenStudy (anonymous):

the two ratios are equivalent when you cross multiply

OpenStudy (anonymous):

equate both ratios

OpenStudy (anonymous):

if both sides are equal then u have proof

OpenStudy (anonymous):

Okay, thanks guys :)

OpenStudy (michele_laino):

I think it is: \[\frac{ 4 }{ 6 }\times \frac{ 21 }{ 14 }=\frac{ 2 }{ 2 }\times \frac{ 7 }{ 7 }=1\]